# Image Generation Models: A Technical History

### 08 Mar 2026 v1

@@ 3.1 Training and Objective Function
They also proposed another approach called minibatch discrimination, and here is how it works. During mode collapse, $G$ emits near-identical samples, and $D$ does not have a way to penalize the lack of diversity. To alleviate this, they give $D$ features that measure the closeness among samples in the same minibatch. Additionally, in order to damp the spiral of two-player gradient dynamics, they nudge the parameters towards the running average of the past values. Arjovsky et al. [29] provide a comprehensive analysis of this problem and prove that the data distribution $p_{\text{data}}$ and the model distribution $p_{g}$ live on different low-dimensional manifolds with nearly disjoint support. As a result, an almost perfect discriminator exists. Their recommendation is to smooth the two distributions with noise so that the discriminator cannot perfectly separate real versus generated data.
@@
Moreover, instead of training a discriminative classifier, WGAN authors train a critic that gives higher scores to real images. The distance between the average real and fake scores is an estimation of the Wasserstein distance — the smaller the gap, the closer the distributions. They also prevent the critic from changing too fast by clamping every weight to a fixed box after each optimization step.
@@
To illustrate the impact of their approach, Fig. 16 shows the non-vanishing gradients across a wider region of parameter space in WGAN compared to the vanishing gradients in regular GANs.
Weight clipping in WGANs can result in undesirable behaviors like generating poor samples or failing to converge. Gulrajanai et al. [31] proposed an alternative, WGAN-GP. Instead of clipping weights, they penalize the norm of gradient of the critic with respect to its input. With this small change, they are able to achieve stable training in a wide variety of GAN architectures with no hyperparameter tuning. Box 4 shows their overall algorithm which is a slight modification of WGAN.
@@ 3.2 Conditional GANs
Shortly after the original GAN paper, Mirza and Osindero [35] introduced the Conditional GANs (CGANs). CGANs can be constructed by feeding the class label $y$ to both the generator and discriminator, as shown in the top section of Fig. 17. The objective function was represented as:
@@
Later on Odena et al. [36] introduced Conditional Image Synthesis with Auxiliary Classifier GANs (AC-GANs) as shown in the bottom section of Fig. 17. They define $x_{\text{fake}}$ as below, where $c$ is the class label and $z$ is the input noise.
@@
Koyama and Miyato [37] modified the prior conditional GAN framework. Instead of concatenating the label $y$ into the discriminator, they define a discriminator score that is the sum of $\langle v_{y},\phi(x)\rangle$ and $\psi(\phi(x))$. This first term measure whether the image belongs to class $y$ and the second terms measures whether the image look real.
@@
where $\phi(x)$ denotes the input image ($x$) features derived from the discriminator network, $v_{y}$ is the embedding vector learned for class $y$, $\psi(\cdot)$ is a small network that outputs a scalar measuring how realistic the image looks. Making this adjustment, they are able to achieve an Inception Score of $29.7\pm 0.61$ compared to $28.5\pm 0.2$ in AC-GANs on the ImageNet. Refer to Table I for more detailed results. Fig. 19 shows how they are able to generate a more realistic photo of a typewriter compared to the previous concatenation methods for labels.
@@
*projection (850K iterations) 36.8 $\pm$ 0.44
TABLE I: Inception Scores on ImageNet
@@ 3.3 High Quality Image Generation
Around 2018 and onward, series of papers were introduced that dramatically increased the quality of generated images and provided more control over the style of the generated output. Karras et al. [38] introduced a new way of training GANs. Their training starts with very low-resolution images of $4\times 4$ pixels. They increase network complexity and image size progressively as training continues. Fig. 20 shows their approach and sample images generated. Following [27], they used the minibatch standard deviation to provide a diversity score for images. Additionally, they incorporated pixel-wise vector normalization in the generator.
Figure 20: Authors start from low-resolution images and progressively increase the complexity of the models and the size of the input images. Image adapted from [38].
@@
In StyleGAN1, Karras and coauthors [39] borrowed ideas from style transfer architectures. Instead of feeding $z$ directly into the first layer, they mapped $z$ to an intermediate latent vector $w$, as shown in Fig. 21. Early layers set the pose and shape in generated images, while later layers determine color and fine details. Another interesting concept they introduced is style mixing, summarized in Box 5.
Figure 21: StyleGAN1 architecture where $z$ is mapped to an intermediate latent $w$. Early layers determine coarse structure, while later layers refine color and details. Sample pictures generated by StyleGAN. Image source: [39]
Fig. 22 shows how they are able to generate high-quality variant of source $A$ images by using images in source $B$ for coarse style.
@@
In StyleGAN3, the authors faced a challenge known as texture sticking, where image details were locked to a specific pixel coordinate as shown in Fig. 24. To address this, the authors borrowed ideas from signal processing to remove aliasing effects. They incorporated anti-aliasing techniques inside the generator. For instance, they avoided input features that could encode pixel coordinates. Additionally, during upsampling or downsampling, they applied anti-aliasing filters. Moreover, when nonlinearities were applied, they increased the internal resolution, applied the nonlinearity, and then filtered back. This ensured that image features did not get stuck at specific pixel coordinates. Using these techniques, they were able to improve both FID and EQ scores compared to StyleGAN2 set forth in Table II. The Fréchet Inception Distance(FID) was introduced by Heusel et al. [42] and it measures how close the distribution of generated images is to the distribution of real images. Equivariance scores, EQ-T [43] and EQ-R [41], measure whether image features stick to the image or to the screen as the image rotates or translates. Ideally we would like the features to stick to the image as they are rotated or translated and not the screen(i.e., fixed pixel location) like in Fig. 24.
@@ 3.4 Other Applications of GANs
where $\Phi(\cdot)$ is a pre-trained VGG [49] network that acts as a feature extractor. $I_{LR}$ and $I_{HR}$ are low- and high-resolution images, respectively. The parameter $\lambda$ balances realism and faithfulness. Larger $\lambda$ corresponds to sharper details but comes with a higher risk of hallucination, while smaller $\lambda$ leads to smoother details but fewer hallucinations.
Using this approach, they are able to recover photo-realistic textures from downsampled images, as shown in Fig. 27.
Figure 27: Image on the left is a 4x upsampled version of the image on the right.
@@ 3.5 Conclusion
Generative Adversarial Networks (GANs) have had a long-lasting impact on image generation. They started in 2014 with low-quality (but still impressive) results, and over the next four to five years, the quality of images improved dramatically. The idea was original and quickly drew attention, but it came with downsides. First, as discussed earlier, they are hard to optimize, and the generator can fall into mode collapse, where many outputs look the same. Second, GAN training can be sensitive to optimizer and hyperparameter choices. Third, with limited data, the discriminator can overfit and the generator will suffer greatly from this. Lastly, metrics such as the Inception Score(IS) and Fréchet Inception Distance (FID) may not reliably correlate with the perceptual quality of the images.
@@ 4 Normalizing Flows
Around 2015, generative image models moved from niche to mainstream research interest. VAEs, GANs, and Normalizing Flows (NF) were introduced within two years of each other. The general idea behind normalizing flows is to transform data to a simple distribution (e.g., Gaussian) via an invertible mapping without losing any information [50, 51]. To achieve this goal, a step by step transformation is learned in order to transform the input into a simple distribution such that each step is reversible. Dinh et al. [52, 53] popularized the idea in the context of generative image modeling.
Dinh et al. [52] pass the input through a transformation such that the resulting output fits a distribution $p_{Z}$ that consists of $d$ independent components, $z_{i}$. The authors’ chosen distribution was Gaussian or logistic.
@@
The transformation $z=f(x)$ is invertible where $x=f^{-1}(z)$. One of the key design choices they make in this paper is the choice of $f$. It must be invertible and the dimension of $z$ is the same as $x$. Fig. 28 shows the simplified diagram of normalizing flows. $f$ and $f^{-1}$ share the same neural network parameters.
Using the change of variable formula leads to (for proof refer to  [54]):
@@
In practice, since the determinant is 1, they add a scale at the end to allow the model to give more weight to some dimensions and less on others. The overall training and generation steps of the NICE [52] paper can be summarized as in Box 6.
@@
Fig. 29 shows unbiased samples from a NICE model. They sample $z\sim p_{Z}(z)$ and output $x=f^{-1}(z)$.
@@
Dinh et al. [53] did a follow-up paper, ”Density Estimation Using Real NVP,” where they improve their previous results [52]. They made a few changes compared to their earlier work (NICE). They switched from additive coupling as shown in Eq. 50 to affine coupling, where
@@ 4.1 Other Variants of Normalizing Flows
Later, Rezende and Mohamed [55] introduced Normalizing Flows in the context of VAEs to improve the expressiveness of the posterior distribution. They used flows to parameterize the posterior in VAEs. They started from a base density $q_{0}(z_{0}|x)$ and passed it through a sequence of invertible maps
@@
Kingma and Dhariwal [59] improve the previous state of the art results in Normalizing Flows and produce high fidelity sample images. They propose three changes to Normalizing Flow architectures.
First, they replace the batch normalization layer with activation normalization (ActNorm). ActNorm performs an affine transformation of activations using a per-channel scale and bias. They set it up such that activations are zero-mean and unit-variance on the first minibatch, afterwards the model learns. This change has a stabilizing impact on deep networks and large images. Second, they replace the fixed permutation used in prior work [53] with a learnable invertible $1\times 1$ convolution layer. Third, they maintain the affine coupling layer introduced by Nice/RealNVP [53, 52], but zero-initialize the last convolution layer of each neural network. They perform splits only along the channel dimension to simplify the overall architecture. Fig. 31 shows a comparison of their architecture compared to RealNVP [53]. By making these changes, they are able to generate realistic images such as those shown in Fig. 32.
@@
Ho, Chen et al. [60] introduced Flow++. They propose three key changes to the flow-based generative models. First, instead of adding uniform dequantization noise to pixels, Flow++ learns a conditional noise model $q(u|x)$ (parameterized by a small flow). They train this auxiliary flow with a tight variational bound, which leads to a better objective and reduces overfitting. Second, they incorporate more expressive coupling layers to better capture data complexity. Third, to improve long-range dependencies, they incorporate convolutional layers and multi-head self-attention inside the network that predicts the coupling parameters. By making these changes, Flow++ achieves state-of-the-art performance among non-autoregressive generative models at the time (2019).
They evaluate using the metric bits per dimension (bits/dim), defined as
$\text{bits/dim}=\frac{\text{NLL}}{N\log(2)}$ (61)
where $N$ is the number of pixels and NLL is the negative log likelihood. Lower bits/dim indicate a better generative model. Table III shows the results of Flow++ compared to other models discussed earlier.
@@
Flow++ (theirs) 3.08 3.86 3.69
TABLE III: Unconditional image modeling results in bits/dim. Table source:  [60].
@@
Prior super-resolution works favor the average of many plausible high-resolution images, as a result the output is blurry. Since SRFlow learns a conditional density $p(y|x)$, it can sample and generate multiple photorealistic images as shown in Fig. 34.
@@ 4.2 Modern Directions in Normalizing Flows
Papamakarios and coauthors [51] published a comprehensive review paper that provides a unified perspective on normalizing flows. They show that normalizing flows can represent any distribution $p_{x}(x)$ under reasonable conditions on $p_{x}(x)$. More specifically, for two well-behaved distributions, $p_{x}(x)$ (target) and $p_{u}(u)$ (base), there exists an invertible function that can turn $p_{u}(u)$ into $p_{x}(x)$. Moreover, they offer a unifying view on how to use normalizing flows for generative modeling versus also density estimation.
@@ 5 Transformer and Autoregressive Models
This section groups together image generation models that make predictions autoregressively regardless of architecture. Early models achieved this with recurrent or convolutional networks, while later architectures adopted transformers [66] and all of them follow the same next token (patch) prediction principle. In autoregressive models, we predict the next element conditioned on everything generated so far. In the case of images, an image is serialized (via pixels or discrete latents) and generation conditions on previously generated tokens.
@@ 5.1 RNN and CNN Autoregressive Image Generation
They introduced two different architectures: PixelRNN and PixelCNN. PixelRNN is an LSTM-based architecture that utilizes masked convolution for computing features and LSTM to model dependencies across previously generated pixels. In order to provide parallelization during training, they also introduced PixelCNN. PixelCNN is a convolutional-based architecture that causally models the dependence among pixels via masked convolution and a stack of convolutional layers. During training, the model factorizes the likelihood in raster-scan order (left-to-right, top-to-bottom) using masked convolutions. While doing this, they use a masking mechanism to make sure that each pixel can see only the past rows and columns. The overall architecture for PixelCNN and LSTM-based PixelRNN is shown in Table IV.
@@
They train by minimizing the negative log-likelihood of a 256-way softmax for each pixel:
@@
Shortly after Oord et al. [9] tried to improve PixelCNN by introducing a gating mechanism. The authors discovered that PixelRNN tended to perform better than PixelCNN because of the gating mechanism in LSTMs. In order to achieve a similar behavior in PixelCNN, they replaced the ReLU block with a gated convolutional block:
@@
Fig. 37 shows sample images generated conditioned on Lhasa Apso dog (top) and Brown bear (bottom).
@@
They also tried using PixelCNN as the decoder in a standard autoencoder. They used conditional PixelCNN for modeling (decoding) $p(x\mid h)$, where $h$ is the latent code generated by the encoder. They achieved promising results in image reconstruction using this approach.
Salimans et al. [68] introduce PixelCNN++, in which they make 5 changes to the original model’s structure and training. Those changes include: 1) instead of a 256-way softmax, they use a discretized logistic mixture model on pixels and round it to the nearest 8-bit value $z\in\{0,\ldots,255\}$. 2) They simplify the masking logic by conditioning the network on previous whole pixels. 3) They utilize down-sampling and up-sampling to capture structure at multiple resolutions. 4) They include long-range skip connections to recover fine details and help with optimization. 5) Apply dropout for regularization. Applying these changes, they are able to improve bits per sub-pixel from 3.00 in PixelCNN to 2.92 in PixelCNN++.
Chen and coauthors [69] introduced PixelSNAIL. They combine causal convolution with self-attention in order to better capture long-range dependencies in autoregressive generation like PixelCNN. One issue with convolutions is that they have a finite receptive field and cannot capture information from distant pixels. They use self-attention to counteract this issue. They interleave causal convolutions and self-attention so that each pixel can read from local context as well as distant pixels. Fig. 38 shows the overall architecture of their model. Using this approach, they are able to improve bits per sub-pixel from 2.92 in PixelCNN++ to 2.85.

### 29 Mar 2026 v2

@@ 3.1 Training and Objective Function
They also proposed another approach called minibatch discrimination where during mode collapse, $G$ emits near-identical samples, and $D$ does not have a way to penalize the lack of diversity. To alleviate this, they give $D$ features that measure the closeness among samples in the same minibatch. Additionally, in order to damp the spiral of two-player gradient dynamics, they nudge the parameters towards the running average of the past values. Arjovsky et al. [29] provide a comprehensive analysis of this problem and prove that the data distribution $p_{\text{data}}$ and the model distribution $p_{g}$ live on different low-dimensional manifolds with nearly disjoint support. As a result, an almost perfect discriminator exists. Their recommendation is to smooth the two distributions with noise so that the discriminator cannot perfectly separate real versus generated data.
@@
Moreover, instead of training a discriminative classifier, WGAN authors train a critic that gives higher scores to real images. The distance between the average real and fake scores is an estimation of the Wasserstein distance, where a smaller gap means closer distributions. They also prevent the critic from changing too fast by clamping every weight to a fixed box after each optimization step. To illustrate the impact of their approach, Fig. 16 shows the non-vanishing gradients across a wider region of parameter space in WGAN compared to the vanishing gradients in regular GANs.
@@
Clamping weights in WGANs can result in undesirable behaviors like generating poor samples or failing to converge. Gulrajani et al. [31] proposed an alternative, WGAN-GP. Instead of clipping weights, they penalize the norm of the gradient of the critic with respect to its input. With this small change, they are able to achieve stable training in a wide variety of GAN architectures with no hyperparameter tuning. Box 4 shows their overall algorithm, which is a slight modification of WGAN.
@@ 3.2 Conditional GANs
Shortly after the original GAN paper, Mirza and Osindero [35] introduced the Conditional GANs (CGANs). CGANs can be constructed by feeding the class label $y$ to both the generator and the discriminator, as shown in the top section of Fig. 17. The objective function was represented as:
@@
Later, Odena et al. [36] introduced Conditional Image Synthesis with Auxiliary Classifier GANs (AC-GANs) as shown in the bottom section of Fig. 17. They define $x_{\text{fake}}$ as below, where $c$ is the class label and $z$ is the input noise.
@@
Koyama and Miyato [37] modified the prior conditional GAN framework. Instead of concatenating the label $y$ into the discriminator, they define a discriminator score that is the sum of $\langle v_{y},\phi(x)\rangle$ and $\psi(\phi(x))$. The first term measures whether the image belongs to class $y$ and the second term measures whether the image looks real.
@@
where $\phi(x)$ denotes the input image ($x$) features derived from the discriminator network, $v_{y}$ is the embedding vector learned for class $y$, and $\psi(\cdot)$ is a small network that outputs a scalar measuring how realistic the image looks. Making this adjustment, they are able to achieve an Inception Score of $29.7\pm 0.61$ compared to $28.5\pm 0.2$ in AC-GANs on the ImageNet. Refer to Table I for more detailed results. Fig. 19 shows how they are able to generate a more realistic photo of a typewriter compared to the previous concatenation methods for labels.
@@
projection (850K iterations) 36.8 $\pm$ 0.44
TABLE I: Comparing the Inception Scores of different conditional GAN models on the ImageNet
@@ 3.3 High Quality Image Generation
From around 2018 onward, a series of papers were introduced that dramatically increased the quality of generated images and provided more control over the style of the generated output. Karras et al. [38] introduced a new way of training GANs. Their training starts with very low-resolution images of $4\times 4$ pixels. They increase network complexity and image size progressively as training continues. Fig. 20 shows their approach and sample images generated. Following [27], they used the minibatch standard deviation to provide a diversity score for images. Additionally, they incorporated pixel-wise vector normalization in the generator.
Figure 20: (Top) Karras et al. [38] start from low-resolution images and progressively increase the complexity of the models and the size of the input images. (Bottom) Sample images generated from this approach. Image adapted from [38].
@@
In StyleGAN1, Karras and coauthors [39] borrowed ideas from style transfer architectures. Instead of feeding $z$ directly into the first layer, they mapped $z$ to an intermediate latent vector $w$, as shown in Fig. 21. Early layers set the pose and shape in generated images, while later layers determine color and fine details. Another interesting concept they introduced is style mixing, summarized in Box 5. Fig. 22 shows how they are able to generate high-quality variant of source AA images by using images in source $B$ for coarse style.
Figure 21: (Top) StyleGAN1 architecture where $z$ is mapped to an intermediate latent $w$. Early layers determine coarse structure, while later layers refine color and details. (Bottom) Sample pictures generated by StyleGAN. Image source: [39]
@@
In StyleGAN3, the authors faced a challenge known as texture sticking, where image details were locked to a specific pixel coordinate as shown in Fig. 24. To address this, the authors borrowed ideas from signal processing to remove aliasing effects. They incorporated anti-aliasing techniques inside the generator. For instance, they avoided input features that could encode pixel coordinates. Additionally, during upsampling or downsampling, they applied anti-aliasing filters. Lastly, when nonlinearities were applied, they first increased the internal resolution, applied the nonlinearity, and then filtered back. This ensured that image features did not get stuck at specific pixel coordinates. Using these techniques, they were able to improve both FID and EQ scores compared to StyleGAN2 set forth in Table II. The Fréchet Inception Distance (FID) was introduced by Heusel et al. [42] and it measures how close the distribution of generated images is to the distribution of real images. Equivariance scores, EQ-T [43] and EQ-R [41], measure whether image features stick to the image or to the screen as the image rotates or translates. Ideally we would like the features to stick to the image as they are rotated or translated and not the screen(i.e., fixed pixel location) like in Fig. 24.
@@ 3.4 Other Applications of GANs
Figure 27: Image on the left is a 4x upsampled version of the image on the right. Image source:  [48]
@@
where $\Phi(\cdot)$ is a pre-trained VGG [49] network that acts as a feature extractor. $I_{LR}$ and $I_{HR}$ are low- and high-resolution images, respectively. The parameter $\lambda$ balances realism and faithfulness. Larger $\lambda$ corresponds to sharper details but comes with a higher risk of hallucination, while smaller $\lambda$ leads to smoother details but fewer hallucinations. Using this approach, they are able to recover photo-realistic textures from downsampled images, as shown in Fig. 27.
@@ 3.5 Conclusion
Generative Adversarial Networks (GANs) have had a long-lasting impact on image generation. They started in 2014 with low-quality (but still impressive) results, and over the next four to five years, the quality of images improved dramatically. The idea was original and quickly drew attention, but it came with downsides. First, as discussed earlier, they are hard to optimize, and the generator can fall into mode collapse, where many outputs look the same. Second, GAN training can be sensitive to optimizer and hyperparameter choices. Lastly, with limited data, the discriminator can overfit and the generator will suffer greatly from this.
@@ 4 Normalizing Flows
Around 2015, generative image models moved from niche to mainstream research interest. VAEs, GANs, and Normalizing Flows (NF) were introduced within two years of each other. The general idea behind normalizing flows is to transform data to a simple distribution (e.g., Gaussian) via an invertible mapping without losing any information [50, 51]. To achieve this goal, a step by step transformation is learned in order to transform the input into a simple distribution such that each step is reversible. Dinh et al. [52, 53] popularized the idea in the context of generative image modeling by passing the input through a transformation such that the resulting output fits a distribution $p_{Z}$ that consists of $d$ independent components, $z_{i}$. The authors’ chosen distribution was Gaussian or logistic.
@@
The transformation $z=f(x)$ is invertible where $x=f^{-1}(z)$. One of the key design choices they make is the choice of $f$. It must be invertible and the dimension of $z$ is the same as $x$. Fig. 28 shows the simplified diagram of normalizing flows. $f$ and $f^{-1}$ share the same neural network parameters. Using the change of variable formula leads to (for proof refer to  [54]):
@@
In practice, since the determinant is 1, they add a scale at the end to allow the model to give more weight to some dimensions and less on others. The overall training and generation steps of the NICE [52] paper can be summarized as in Box 6. Fig. 29 shows unbiased samples from a NICE model. They sample $z\sim p_{Z}(z)$ and output $x=f^{-1}(z)$.
@@
Dinh et al. [53] published a follow-up paper, ”Density Estimation Using Real NVP,” where they improve their previous results [52]. They made a few changes compared to their earlier work (NICE). They switched from additive coupling as shown in Eq. 50 to affine coupling, where
@@ 4.1 Other Variants of Normalizing Flows
Rezende and Mohamed [55] introduced Normalizing Flows in the context of VAEs to improve the expressiveness of the posterior distribution. They used flows to parameterize the posterior in VAEs. They started from a base density $q_{0}(z_{0}|x)$ and passed it through a sequence of invertible maps
@@
Later, Kingma and Dhariwal [59] improve the previous state of the art results in Normalizing Flows and produce high fidelity sample images. They propose three changes to Normalizing Flow architectures. First, they replace the batch normalization layer with activation normalization (ActNorm). ActNorm performs an affine transformation of activations using a per-channel scale and bias. They set it up such that activations are zero-mean and unit-variance on the first minibatch, afterwards the model learns. This change has a stabilizing impact on deep networks and large images. Second, they replace the fixed permutation used in prior work [53] with a learnable invertible $1\times 1$ convolution layer. Third, they maintain the affine coupling layer introduced by Nice/RealNVP [53, 52], but zero-initialize the last convolution layer of each neural network. They perform splits only along the channel dimension to simplify the overall architecture. Fig. 31 shows a comparison of their architecture compared to RealNVP [53]. By making these changes, they are able to generate realistic images such as those shown in Fig. 32.
@@
Ho, Chen et al. [60] introduced Flow++. They propose three key changes to the flow-based generative models. First, instead of adding uniform dequantization noise to pixels, Flow++ learns a conditional noise model $q(u|x)$ (parameterized by a small flow). They train this auxiliary flow with a tight variational bound, which leads to a better objective and reduces overfitting. Second, they incorporate more expressive coupling layers to better capture data complexity. Third, to improve long-range dependencies, they incorporate convolutional layers and multi-head self-attention inside the network that predicts the coupling parameters.
@@
Flow++ 3.08 3.86 3.69
TABLE III: Unconditional image modeling results in bits/dim. Table source [60].
By making these changes, Flow++ achieves state-of-the-art performance among non-autoregressive generative models at the time (2019). They evaluate using bits per dimension (bits/dim), defined as
$\text{bits/dim}=\frac{\text{NLL}}{N\log(2)}$ (61)
where NN is the number of pixels and NLL is the negative log likelihood. Lower bits/dim indicate a better generative model. Table III shows the results of Flow++ compared to other models discussed earlier.
@@
Prior super-resolution works favor the average of many plausible high-resolution images, as a result the output is blurry. However, SRFlow learns a conditional density $p(y|x)$, as a result it can sample and generate multiple photorealistic images as shown in Fig. 34.
@@ 4.2 Modern Directions in Normalizing Flows
Papamakarios and coauthors [51] published a comprehensive review paper that provides a unified perspective on normalizing flows. They show that normalizing flows can represent any distribution $p_{x}(x)$ under reasonable conditions on $p_{x}(x)$. More specifically, for two well-behaved distributions, $p_{x}(x)$ (target) and $p_{u}(u)$ (base), there exists an invertible function that can turn $p_{u}(u)$ into $p_{x}(x)$. Moreover, they offer a unifying view on how to use normalizing flows for generative modeling versus density estimation.
@@ 5 Transformer and Autoregressive Models
This section focuses on image generation models that make predictions autoregressively. Early models achieved this with recurrent or convolutional networks, while later architectures adopted transformers [66] and all of them follow the same next token (patch) prediction principle. In autoregressive models, we predict the next element conditioned on everything generated so far. In the case of images, an image is tokenized (via pixels or discrete latents) and generation conditions on previously generated tokens.
@@ 5.1 RNN and CNN Autoregressive Image Generation
They introduced two different architectures: PixelRNN and PixelCNN. PixelRNN is an LSTM-based architecture that utilizes masked convolution for computing features and LSTM to model dependencies across previously generated pixels. In order to provide parallelization during training, they also introduced PixelCNN. PixelCNN is a convolutional-based architecture that causally models the dependence among pixels via masked convolution and a stack of convolutional layers. During training, the model factorizes the likelihood in raster-scan order (left-to-right, top-to-bottom) using masked convolutions. While doing this, they use a masking mechanism to make sure that each pixel can see only the past rows and columns. The overall architecture for PixelCNN and LSTM-based PixelRNN is shown in Table IV. They train by minimizing the negative log-likelihood of a 256-way softmax for each pixel:
@@
Shortly after, Oord et al. [9] tried to improve PixelCNN by introducing a gating mechanism. The authors discovered that PixelRNN tended to perform better than PixelCNN because of the gating mechanism in LSTMs. In order to achieve a similar behavior in PixelCNN, they replaced the ReLU block with a gated convolutional block:
@@
Fig. 37 shows sample images generated conditioned on Lhasa Apso dog (top) and Brown bear (bottom). They also tried using PixelCNN as the decoder in a standard autoencoder. They used conditional PixelCNN for modeling (decoding) $p(x\mid h)$, where $h$ is the latent code generated by the encoder.
@@
Salimans et al. [68] introduced PixelCNN++, in which they made 5 changes to the original PixelCNN’s model structure and training. Those changes include: 1) instead of a 256-way softmax, they use a discretized logistic mixture model on pixels and round it to the nearest 8-bit value $z\in\{0,\ldots,255\}$. 2) They simplify the masking logic by conditioning the network on previous whole pixels. 3) They utilize down-sampling and up-sampling to capture structure at multiple resolutions. 4) They include long-range skip connections to recover fine details and help with optimization. 5) Apply dropout for regularization. By applying these changes, they are able to improve bits per sub-pixel from 3.00 in PixelCNN to 2.92 in PixelCNN++.
Chen and coauthors [69] introduced PixelSNAIL. They combine causal convolution with self-attention in order to better capture long-range dependencies in autoregressive generation models like PixelCNN. One issue with convolutions is that they have a finite receptive field and cannot capture information from distant pixels. They use self-attention to counteract this issue. They interleave causal convolutions and self-attention so that each pixel can read from local context as well as distant pixels. Fig. 38 shows the overall architecture of their model. Using this approach, they are able to improve bits per sub-pixel from 2.92 in PixelCNN++ to 2.85.